OpenStudy (anonymous):

Factoring Question: Factor the following a)8a^2-2ab-21b^2

OpenStudy (anonymous):

(2a+3b)(4a-7b)

OpenStudy (anonymous):

How did you do that ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

on sec ..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

If you can't factor this kind of quadratic, omit the second variable (but remember to put it back at the end) and proceed the same way as above 8a^2 - 2a - 21 = 0 a = -3/2, 7/4 (a + 3/2) (a - 7/4) 8 (a + 3/2) (a - 7/4) (2) (4) (a + 3/2) (a - 7/4) (2a + 3) (4a - 7) *(2a + 3b) (4a - 7b)*

OpenStudy (anonymous):

How are you getting -3 and 7

OpenStudy (anonymous):

I was taught to multiply the a by the c

OpenStudy (anonymous):

then find two numbers that add to the "b" and multiply to c and a together

OpenStudy (anonymous):

well i was taught like this either way the answer is the same :)

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

if we do it my way what two numbers multiply to -168 and add to -2?

OpenStudy (anonymous):

i can't figure that part out

OpenStudy (anonymous):

-14*12=-168 -14+12=-2

OpenStudy (anonymous):

Oh, how did you find that so fast. I can't figure out numbers when they are so inlarged

OpenStudy (anonymous):

lol you just have to pratice and study your times tables :)

OpenStudy (anonymous):

aha yeah truee, thanks again very much i appreciate the help

OpenStudy (anonymous):

You Are Very Welcome :)

OpenStudy (anonymous):

Do you mind trying to do another factoring equation that is a bit different?

OpenStudy (anonymous):

If you are busy it is okay.

OpenStudy (anonymous):

Yah im about to leave some where sorry

OpenStudy (anonymous):

Okay thats alright good bye

OpenStudy (anonymous):

bye but thanks for the medal

OpenStudy (anonymous):

No thank you for the help :)